# From Spin Waves to Monte Carlo Simulations: Compiling an Experimental Exchange Interaction Dataset for Magnetic Materials

### v1

@@ Abstract
Inelastic neutron scattering data on magnetic crystals are highly valuable in materials science, as they provide direct insight into microscopic magnetic interactions. Using spin wave theory, these interactions can be extracted from magnetic excitations observed in such experiments. However, these datasets are often scattered across the literature and lack standardization, limiting their accessibility and usability. In this paper, we compile and standardize Heisenberg exchange interaction data for magnetic materials obtained from inelastic neutron scattering experiments. Through an extensive literature review, we identify experimental data for approximately 100 magnetic materials. The standardized dataset includes mapping the results of various Heisenberg Hamiltonians into a unified standard form, visualizations of crystal structures with annotated exchange interactions, and input and output files from Monte Carlo simulations performed for each compound using the ESpinS code. Using experimentally determined exchange interactions, we calculate transition temperatures ($T_{c}$) via classical Monte Carlo simulations. Additionally, we assess the effectiveness of the $(S+1)/S$ correction within classical Monte Carlo simulations, finding that it produces transition temperatures in excellent agreement with experimental values in most cases. The complete dataset, along with supporting resources, is publicly available on GitHub.
@@ I Introduction
Despite the availability of theoretical tools such as LSWT for analyzing INS data, INS measurements remain limited due to the specialized and resource-intensive equipment required. Our literature review identifies only about 100 studies that report both magnon INS data and corresponding spin model mappings using LSWT. Compiling and standardizing this data into a unified framework would greatly benefit magnetism research.
@@
Building on our previous findings, this work aims to systematically assess the influence of the $(S+1)/S$ correction on magnetic transition temperatures across a broader range of materials. We extend our dataset to encompass approximately 73 inelastic neutron scattering (INS) studies, ensuring that all extracted exchange parameters are standardized within a consistent spin model Hamiltonian. Using classical MC simulations, we compute transition temperatures both with and without the correction. Our analysis confirms that incorporating the $(S+1)/S$ factor significantly improves the agreement between simulated and experimental $T_{c}$ values. To support further research and reproducibility, the complete dataset has been made publicly available on GitHub [11].
@@ II.1 Data gathering and standardizing
Figure 1: Illustrative images of sample crystal structures from our database. Each image highlights the exchange interactions and their labeling, as presented in the corresponding INS paper. The labels in parentheses indicate the ranking of exchange interactions based on nearest-neighbor distances (e.g., $J_{1}$ corresponds to the first nearest neighbor,$J_{2}$ to the second nearest neighbor, and so on).
One of the key challenges in standardizing exchange interactions is the ambiguity in the choice of spin Hamiltonian conventions, particularly regarding the double counting of pairwise interactions. For example, models that use $-\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $-\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ are designed so that each interaction between spins at sites $i$ and $j$ is counted only once. In contrast, models that adopt $-\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $-2\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ count each pairwise interaction twice. A major source of confusion arises when some studies claim to use the convention $-\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, yet a detailed analysis—often involving comparisons with other works—reveals that their reported exchange constants have already been corrected to avoid double counting. In practice, these studies effectively use the convention $-\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, as seen, for example, in Refs. [12, 13, 14, 15, 3].
@@ II.3 Data on Github
III Results and discussion
III.1 Predicting transition temperature using Monte Carlo
@@
Figure 2 shows the transition temperatures obtained from MC simulations ($T_{\mathrm{MC}}$) for all 73 compounds, revealing a significant deviation from the experimental transition temperatures. However, applying the $(S+1)/S$ correction:
@@
The MAPE for $T_{\mathrm{MC}}^{*}$ is 9.6%, whereas for $T_{\mathrm{MC}}$, the error is much greater, equal to 35.1%. For 37% of the compounds, the absolute percentage error (APE) is less than 5%. In 30% of the cases, APE falls between 5% and 10%, while for 12% of the compounds, it ranges from 10% to 15%. Only 20% of the compounds exhibit an APE greater than 15%. The detailed data from the MC simulations shown in Fig. 2 are provided in Appendix A.
It is important to note that exchange interactions derived from INS data are subject to uncertainties due to their reliance on pure fitting to LSWT. Additionally, exchange interactions can vary with temperature in practice [31, 32, 33]. These factors collectively impact the accuracy of predicting the transition temperature using simulation methods such as MC.
@@
III.2 Refining Spin Hamiltonians Using Monte Carlo Results
@@
Figure 2: Results from MC simulations for 73 compounds. $T_{\mathrm{MC}}$ represents the transition temperature obtained from MC simulations using exchange parameters derived from INS data fitted with spin-wave theory. $T_{\mathrm{MC}}^{*}$ denotes the quantum-corrected transition temperature, obtained by applying the correction factor $(S+1)/S$ to $T_{\mathrm{MC}}$. All data points shown in this diagram are provided in Appendix A.
@@
[…]
@@
[…]
@@
During our investigation, we found instances where the transition temperature obtained from Monte Carlo simulations ($T_{\mathrm{MC}}$) exceeded the experimental value, and other cases where the quantum-corrected temperature ($T_{\mathrm{MC}}^{*}$) was significantly lower. Such discrepancies often indicate problems with the INS data or the fitting procedure. These issues may stem from an incorrectly defined Hamiltonian—such as a missing factor of $1/2$—poor data fitting, or an inaccurate estimate of the spin quantum number $S$. For example, using the INS-derived exchange interactions for LiNiPO4 [36], LiCoPO4 [37], and LiMnPO4 [38], the MC-calculated transition temperatures ($T_{\mathrm{MC}}$) are overestimated. This points to a mismatch between the spin Hamiltonian and the spin-wave formalism, likely caused by an omitted factor of $1/2$ in the Hamiltonian. A closer examination of the spin-wave equations used in these studies supports this conclusion.
IV Conclusions
@@
Through analysis of more than 100 INS studies that derived exchange interactions from spin-wave theory, we identified 73 compounds with reliable data and compiled them into a standardized database. This database includes exchange interaction parameters, spin quantum numbers, reference sources, MC simulation results, structure files, and visualized structures. Using this dataset, we investigated the accuracy of transition temperature predictions by applying the $(S+1)/S$ correction to classical MC results. Our findings show that, in most cases, the corrected results align more closely with experimental values, achieving a mean absolute percentage error (MAPE) of 9%. We propose classical MC simulations and the $(S+1)/S$ correction as useful tools for refining INS data and identifying potential inconsistencies.
@@ Appendix A Transition temperature
$T_{\mathrm{Exp.}}$(K) $T_{\mathrm{MC}}$(K) $T_{\mathrm{MC}}^{*}$(K) S Error(%)
Ba2NiWO6[39] $48$ [39] $23$ $46$ $1$ $4.2$
BaMn2As2[40] $625$[40] $460$ $644$ $5/2$ $3.0$
BaMn2Bi2[41] $387.2$[41] $238$ $333.2$ $5/2$ $13.9$
BaNi2As2O8[42] $18.5$[43] $11$ $22$ $1$ $18.9$
Bi2CuO4[44] $50$ [44] $17.5$ $52.5$ $1/2$ $5$
BiFeO3[45] $650$[45] $431$ $603.4$ $5/2$ $7.2$
Ca3Ru2O7[46] $56$[46] $30$ $60$ $1$ $7.1$
CaMn2Sb2[47] $85$[47] $69$ $115$ $3/2$ $35.3$
CaMn7O12[48] $90$[48] $61$ $122$ $1$ $35.5$
CaMnBi2[49] $270$[49] $209$ $292.6$ $5/2$ $8.4$
CoO[50] $289$[50] $201$ $353$/2323/23 / 2 $15.9$
CoPS3[51] $120$[51] $61$ $101.7$ $3/2$ $15.3$
Cr2O3[52] $308$[52] $194$ $23.33$/2323/23 / 2 $5.0$
Cr2TeO6[53] $93$[53] $53$ $88.3$ $3/2$ $5.0$
Cr2WO6[54] $45$[54] $27$ $45$ $3/2$ $0.0$
CrBr3[55] $32$[55] $20$ $3.33$/2323/23 / 2 $4.2$
CrCl2[56] $20$[56] $15$ $22.5$ $2$ $12.5$
CrCl3[57] $14$[57] $10$ $16.6$ $3/2$ $18.6$
CrI3[58] $61$[58] $39$ $65$ $3/2$ $6.6$
CuFeO2[59] $11$[59] $7.5$ $10.5$ $5/2$ $4.5$
CuO[60] $213$[60] $72$ $216$ $1/2$ $1.4$
EuO[61] $69.15$[61] $52$ $66.9$ $7/2$ $3.3$
EuS[62] $16.57$[62] $13$ $16.7$ $7/2$ $0.8$
Fe2O3[63] $960$[63] $730$ $1022$ $5/2$ $6.4$
FeCl2[64] $23.55$[64] $11.5$ $23$ $1$ $2.3$
FeF2[65] $78.4$[65] $55$ $82.5$ $2$ $5.2$
FeO[66] $198$[66] $148$ $222$ $2$ $12.1$
FePS3[67] $120$[67] $91$ $136.5$ $2$ $13.8$
FePSe3[68] $110$[68] $75$ $112.5$ $2$ $2.3$
FePt3[69] $170$[69] $104$ $167.0$ $1.65$ $1.8$
HoFeO3[70] $647$[70] $456$ $638.4$ $5/2$ $1.3$
KCoF3[71] $114$[72] $33$ $99$ $1/2$ $13.2$
KCuF3[73] $39$[73] $15$ $45$ $1/2$ $15.4$
La2CoO4[74] $275$[74] $161$ $268.3$ $3/2$ $2.4$
La2NiO4[75] $328$[75] $220$ $440$ $1$ $34.1$
LaFeO3[76] $738$[77] $504$ $705.6$ $5/2$ $4.4$
$T_{\mathrm{Exp.}}$(K) $T_{\mathrm{MC}}$(K) $T_{\mathrm{MC}}^{*}$(K) S Error(%)
LaMnO3[78] $139.5$[78] $101$ $151.5$ $2$ $8.6$
LaVO3[76] $150$[79] $70$ $140$ $1$ $6.7$
LiFePO4[80] $50$[80] $41$ $61.5$ $2$ $23$
LuMnO3[81] $87.5$[81] $71$ $106.5$ $2$ $21.7$
MnBi[82] $630$[82] $394$ $591$ $2$ $6.2$
MnBi2Te4[35] $24$[35] $16$ $22.4$ $5/2$ $6.7$
MnF2[16] $67$[16] $48$ $67.2$ $5/2$ $0.3$
MnO[83] $117$[83] $79$ $110.6$ $5/2$ $5.5$
MnO2[84] $92$[84] $63$ $105$ $3/2$ $14.1$
MnPS3[85] $78$[85] $61$ $85.4$ $5/2$ $9.5$
MnPSe3[86] $74$[86] $52$ $72.8$ $5/2$ $1.6$
MnS[87] $152$[87] $105$ $147$ $5/2$ $3.3$
MnS2[88] $48$[88] $25$ $35$ $5/2$ $27.1$
MnTe[89] $310$[89] $251$ $351.4$ $5/2$ $13.4$
MnTiO3[90] $65$[90] $45$ $63$ $5/2$ $3.1$
MnWO4[91] $13.5$[91] $9$ $12.6$ $5/2$ $6.7$
NdFeO3[92] $693$[92] $463$ $648.2$ $5/2$ $6.5$
NiBr2[93] $52$[93] $28$ $56$ $1$ $7.7$
NiCl2[94] $52.3$[94] $36$ $72$ $1$ $37.7$
NiF2[95] $73.2$[95] $42$ $84$ $1$ $14.8$
NiO[34] $523$[34] $310$ $620$ $1$ $18.5$
PrFeO3[96] $707$[96] $471$ $659.4$ $5/2$ $6.7$
Rb2MnF4[97] $38.4$[97] $33$ $46.2$ $5/2$ $20.3$
RbMnCl3[98] $94$[98] $69$ $96.6$ $5/2$ $2.8$
RbMnF3[99] $82.6$[99] $54$ $75.6$ $5/2$ $8.5$
RbNiF3[100] $133$[100] $79$ $158$ $1$ $18.8$
Sr2NiWO6[101] $54$[101] $26$ $52$ $1$ $3.7$
SrMnBi2[102] $290$[102] $219$ $306.6$ $5/2$ $5.7$
SrMnSb2[103] $297$[104] $189$ $283.5$ $2$ $4.5$
TbFeO3[105] $681$[105] $456$ $638.4$ $5/2$ $6.2$
TmFeO3[106] $635$[107] $442$ $618.8$ $5/2$ $2.6$
YbFeO3[108] $627$[109] $440$ $616$ $5/2$ $1.8$
YbMnBi2[13] $290$[13] $203$ $284.2$ $5/2$ $2.0$
YbMnSb2[12] $345$[12] $266$ $372.4$ $5/2$ $7.9$
YFeO3[110] $644.5$[111] $445$ $623$ $5/2$ $3.3$
YMnO3[112] $70$[112] $40$ $60$ $2$ $14.3$
YVO3[113] $118$[114] $59$ $118$ $1$ $0.0$

### v2

# Experimental Exchange Interaction Dataset for Magnetic Materials: Spin Waves to MC Simulations

@@ Abstract
Inelastic neutron scattering (INS) provides direct insights into microscopic magnetic interactions in crystalline materials, making it a valuable experimental technique in condensed matter physics and materials science. These interactions can be extracted by fitting spin wave dispersions to Heisenberg Hamiltonians using spin wave theory. However, such datasets are scattered across the literature and lack a standardized format, which limits their accessibility, reproducibility, and utility. In this work, we compile and standardize exchange interaction data obtained from INS experiments on nearly 100 magnetic materials. The resulting dataset includes exchange parameters expressed in a unified Heisenberg model format, visualizations of crystal structures with annotated exchange pathways, and Monte Carlo simulation files generated using the ESpinS code. We use these experimentally derived exchange interactions to compute magnetic transition temperatures ($T_{c}$) via classical Monte Carlo simulations. Furthermore, we examine the impact of the $(S+1)/S$ correction in the simulations and find it improves agreement with experimental $T_{c}$ values in most cases. All data and related resources are openly available through a public GitHub repository.
@@ I Introduction
Despite the availability of theoretical frameworks such as LSWT for analyzing INS data, the number of reported exchange interactions derived from INS remains limited. This is due to the specialized instrumentation, high-quality large single crystals, and significant beam time required [1, 2, 3], making INS far less common than elastic neutron diffraction for determining magnetic structures  [7, 8, 9]. Our literature survey identified only about 100 studies that report both magnon spectra from INS and corresponding spin-model parameterizations using LSWT. Establishing a unified and standardized database for such results would be a valuable resource for the magnetism community, complementing existing magnetic structure repositories.
@@
Building on our previous findings, this work aims to systematically assess the influence of the $(S+1)/S$ correction on magnetic transition temperatures across a broader range of materials. We extend our dataset to encompass 72 inelastic neutron scattering (INS) studies, ensuring that all extracted exchange parameters are standardized within a consistent spin model Hamiltonian. Using classical MC simulations, we compute transition temperatures both with and without the correction. Our analysis confirms that incorporating the $(S+1)/S$ factor significantly improves the agreement between simulated and experimental $T_{c}$ values. To support further research and reproducibility, the complete dataset has been made publicly available on GitHub [14]. The data is openly accessible and can be freely forked, extended, or updated with new contributions. In addition, for long-term preservation and citability, the dataset has been archived in the Zenodo repository [15].
@@ II.1 Data gathering and standardizing
[…]
One of the key challenges in standardizing exchange interactions is the ambiguity in the choice of spin Hamiltonian conventions, particularly regarding the double counting of pairwise interactions. For example, models that use $\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ are designed so that each interaction between spins at sites $i$ and $j$ is counted only once. In contrast, models that adopt $\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ or $2\sum_{\langle i,j\rangle}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$ count each pairwise interaction twice. A major source of confusion arises when some studies claim to use the convention $\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, yet a detailed analysis—often involving comparisons with other works—reveals that their reported exchange constants have already been corrected to avoid double counting. In practice, these studies effectively use the convention $\frac{1}{2}\sum_{i,j}J_{i,j}\,\mathbf{S}_{i}\cdot\mathbf{S}_{j}$, as seen, for example, in Refs. [16, 17, 18, 19, 3].
@@ III Data Records
IV Technical Validation
IV.1 Predicting transition temperature using Monte Carlo
@@
Figure 2 shows the transition temperatures obtained from MC simulations ($T_{\mathrm{MC}}$) for all 72 compounds, revealing a significant deviation from the experimental transition temperatures. However, applying the $(S+1)/S$ correction:
@@
The MAPE for $T_{\mathrm{MC}}^{*}$ is 9.6%, whereas for $T_{\mathrm{MC}}$, the error is much greater, equal to 35.1%. For 36% of the compounds, the absolute percentage error (APE) is less than 5%. In 32% of the cases, APE falls between 5% and 10%, while for 11% of the compounds, it ranges from 10% to 15%. Only 21% of the compounds exhibit an APE greater than 15%. The detailed data from the MC simulations shown in Fig. 2 are provided in Appendix A.
It is important to note that exchange interactions derived from INS data are subject to several approximations. First, LSWT itself is an approximation to spin-wave theory when fitting INS data to a spin Hamiltonian. Second, in practice, only the dominant exchange terms are typically included in the spin Hamiltonian, while weaker interactions (such as single-ion anisotropy or longer-range couplings) are often neglected. Finally, exchange parameters are not strictly constant but may vary with temperature in practice [35, 36, 37]. Another source of error arises from the classical nature of Monte Carlo simulations: while classical spin models become exact in the large-S limit, quantum effects can play a role for small spin values. Nevertheless, previous studies have shown that for $S\geq 1$ the difference between quantum and classical Monte Carlo results is usually minor, and even for $S=\frac{1}{2}$ the deviations remain small [12, 13]. Finite-size effects in the simulations can also lead to systematic shifts in the transition temperature. Taken together, these experimental and theoretical limitations account for the observed deviations between our predicted and experimental transition temperatures.
@@
IV.2 Refining Spin Hamiltonians Using Monte Carlo Results
@@
Figure 2: Results from MC simulations for 72 compounds. $T_{\mathrm{MC}}$ represents the transition temperature obtained from MC simulations using exchange parameters derived from INS data fitted with spin-wave theory. $T_{\mathrm{MC}}^{*}$ denotes the quantum-corrected transition temperature, obtained by applying the correction factor $(S+1)/S$ to $T_{\mathrm{MC}}$. All data points shown in this diagram are provided in Appendix A.
@@
Figure 3: Magnetic specific heat calculated from MC simulations for three different spin models of MnBi2Te4, based on parameters obtained from INS data. The $J_{1}$$J_{2}$$J_{4}$ model was introduced in Ref. [18], while the $J_{1}$$J_{7}$ and $J_{c}^{\mathrm{aniso}}$ models were proposed in Ref. [39]. The peak of the specific heat defines the MC transition temperature, $T_{\mathrm{MC}}$. The values of $T_{\mathrm{MC}}$ for the $J_{1}$$J_{2}$$J_{4}$, $J_{1}$$J_{7}$, and $J_{c}^{\mathrm{aniso}}$ models are 14.7 K, 15.3 K, and 16.7 K, respectively.
@@
In the first study [18], the authors introduced the $J_{1}$$J_{2}$$J_{4}$ model, which resulted in a MC transition temperature $T_{\mathrm{MC}}=14.7$ K and a corrected transition temperature $T_{\mathrm{MC}}^{*}=20.58$ K (See Fig. 3). In the more recent work [39], the model was extended to include exchange interactions up to the 7th nearest neighbor ($J_{1}$$J_{7}$), and an alternative model incorporating anisotropic inter-layer exchange ($J_{c}^{\mathrm{aniso}}$) was also proposed.
MC simulations using the $J_{1}$$J_{7}$ model yield $T_{\mathrm{MC}}=15.3$ K and $T_{\mathrm{MC}}^{*}=21.42$ K (See Fig. 3). For the $J_{c}^{\mathrm{aniso}}$ model, the results improve further to $T_{\mathrm{MC}}=16.7$ K and $T_{\mathrm{MC}}^{*}=23.38$ K, suggesting that the $J_{c}^{\mathrm{aniso}}$ model provides the best agreement with the experimental data.
@@
During our investigation, we found instances where the transition temperature obtained from Monte Carlo simulations ($T_{\mathrm{MC}}$) exceeded the experimental value, and other cases where the quantum-corrected temperature ($T_{\mathrm{MC}}^{*}$) was significantly lower. Such discrepancies often indicate problems with the INS data or the fitting procedure. These issues may stem from an incorrectly defined Hamiltonian—such as a missing factor of 1/21/2—poor data fitting, or an inaccurate estimate of the spin quantum number $S$. For example, using the INS-derived exchange interactions for LiNiPO4 [40], LiCoPO4 [41], and LiMnPO4 [42], the MC-calculated transition temperatures ($T_{\mathrm{MC}}$) are overestimated. This points to a mismatch between the spin Hamiltonian and the spin-wave formalism, likely caused by an omitted factor of $1/2$ in the Hamiltonian. A closer examination of the spin-wave equations used in these studies supports this conclusion.
@@
Through analysis of more than 100 INS studies that derived exchange interactions from spin-wave theory, we identified 72 compounds with reliable data and compiled them into a standardized database. This database includes exchange interaction parameters, spin quantum numbers, reference sources, MC simulation results, structure files, and visualized structures. Using this dataset, we investigated the accuracy of transition temperature predictions by applying the $(S+1)/S$ correction to classical MC results. Our findings show that, in most cases, the corrected results align more closely with experimental values, achieving a mean absolute percentage error (MAPE) of 9%. We propose classical MC simulations and the $(S+1)/S$ correction as useful tools for refining INS data and identifying potential inconsistencies.
IV.3 Comparison with theories
To provide further insight into the data, we analyze the role of the local atomic environment in determining the exchange interactions. For this purpose, we focus on the bond angle (M1–L–M2) between two magnetic atoms (M1, M2) bridged by a ligand (L), together with the distance between the magnetic atoms (M1–M2). To ensure that the bond angle is well defined, we restrict the analysis to cases where the two magnetic atoms are symmetrically coordinated by the ligand (M1–L = M2–L). Exchange pathways involving more than one intermediate atom (e.g., M1–L–M2–L–M3 or M1–L1–L2–M2) are excluded from the analysis, since the bond angle cannot be uniquely defined in such cases. Figure  4 presents the resulting diagram of exchange interactions as a function of bond angle and interatomic distance. The overall trend is broadly consistent with one of the Goodenough–Kanamori–Anderson (GKA) rules [43, 44, 45]: bond angles close to $180^{\circ}$ generally favor antiferromagnetic interactions ($J<0$), while bond angles near $90^{\circ}$ more often lead to ferromagnetic interactions ($J>0$). It is also important to note that the GKA rules further emphasize the dependence of exchange on the d-orbital filling of the magnetic ions. However, INS data, together with theoretical calculations, clearly demonstrate that the GKA rules are not universally valid. A representative example is KCuF3, where two Cu–F–Cu pathways both have bond angles of $\sim 180^{\circ}$ (corresponding to Cu–Cu distances of 3.92 Å and 4.14 Å), yet one interaction is antiferromagnetic while the other is ferromagnetic . This shows that while the GKA rules provide valuable qualitative guidance, the actual exchange interactions also depend sensitively on factors such as orbital hybridization, covalency, and details of the crystal structure.
Figure 4: The diagram illustrates the exchange interaction as a function of the bond angle (M1–L–M2) between magnetic atoms (M1, M2) and the ligand (L), and the interatomic distance (M1–M2). Red and blue circles represent ferromagnetic and antiferromagnetic interactions, respectively. The circle radius is proportional to the square root of the interaction strength, i.e., $\sqrt{|J|}$.
Comparisons between INS-derived exchange parameters and those obtained from density functional theory (DFT) are highly valuable, but they must be treated with care, as DFT results can vary significantly depending on the choice of basis set, exchange–correlation functional, and other methodological details. To ensure consistency and avoid ambiguities that might arise from mixing results obtained with different approaches (e.g., those available in public databases), we restrict our comparison to two of our previous systematic studies, in which a uniform computational strategy was applied across all compounds. In the first study [46], we employed DFT+$U$ [47] with the Hubbard parameter determined self-consistently using linear response theory, while in the second study [48] we used the meta-GGA r2SCAN [49] functional to calculate exchange interactions. Thirteen compounds are common to both of these earlier works and the present INS dataset. For these systems, we compare the maximum exchange interaction obtained from INS, DFT+$U$, and meta-GGA r2SCAN (Figure  5). The results reveal a clear trend: DFT+$U$ tends to underestimate the exchange interaction strength relative to INS, whereas meta-GGA r2SCAN generally overestimates it. These comparisons, however, should be interpreted with caution. In our earlier DFT studies, the mapping was carried out onto a classical Heisenberg Hamiltonian, while in the present INS analysis the mapping is performed within LSWT, which explicitly incorporates quantum effects. This methodological distinction likely contributes to the observed discrepancies between theory and experiment.
Figure 5: Maximum exchange interaction for 13 compounds, as determined from INS measurements, compared with values obtained from DFT+$U$ [46] and meta-GGA r2SCAN [48].
V Author Contributions
Mojtaba Alaei conducted and supervised the project, performed part of the Monte Carlo (MC) simulations, prepared data for the GitHub repository, and wrote the initial draft of the manuscript. Zahra Mosleh gathered data, prepared inputs for the MC simulations, and contributed to structural visualization. Nafise Rezaei carried out MC simulations, contributed to structural visualization, and participated in manuscript writing. Artem R. Oganov provided intellectual input and financial support. All authors reviewed and edited the manuscript.
VI Competing interests
The authors declare no competing interests
VII Code availability
This study did not involve the development of new code.
@@ Appendix A Transition temperature
$T_{\mathrm{Exp.}}$(K)
$T_{\mathrm{MC}}$(K)
$T_{\mathrm{MC}}^{*}$(K)
S
Error(%)
Ba2NiWO6[50] $48$ [50] $23$ $46$ 11 $4.2$
BaMn2As2[51] 625625[51] $460$ $644$ $5/2$ $3.0$
BaMn2Bi2[52] 387.2387.2[52] $238$ $333.2$ $5/2$ $13.9$
BaNi2As2O8[53] 18.518.5[54] $11$ $22$ 11 $18.9$
Bi2CuO4[55] $50$ [55] $17.5$ $52.5$ $1/2$ 55
BiFeO3[56] 650650[56] $431$ $603.4$ $5/2$ $7.2$
Ca3Ru2O7[57] 5656[57] $30$ $60$ 11 $7.1$
CaMn2Sb2[58] 8585[58] $69$ $115$ $3/2$ $35.3$
CaMn7O12[59] 9090[59] $61$ $122$ 11 $35.5$
CaMnBi2[60] 270270[60] $209$ $292.6$ $5/2$ $8.4$
CoO[61] 289289[61] $201$ $335$ $3/2$ $15.9$
CoPS3[62] 120120[62] $61$ $101.7$ $3/2$ $15.3$
Cr2O3[63] 308308[63] $194$ $323.3$ $3/2$ $5.0$
Cr2TeO6[64] 9393[64] $53$ $88.3$ $3/2$ $5.0$
Cr2WO6[65] 4545[65] $27$ $45$ $3/2$ $0.0$
CrBr3[66] 3232[66] 2020 $33.3$ $3/2$ $4.2$
CrCl2[67] 2020[67] $15$ $22.5$ 22 $12.5$
CrCl3[68] 1414[68] $10$ $16.6$ $3/2$ $18.6$
CrI3[69] 6161[69] $39$ $65$ $3/2$ $6.6$
CuFeO2[70] 1111[70] $7.5$ $10.5$ $5/2$ $4.5$
CuO[71] 213213[71] $72$ $216$ $1/2$ $1.4$
EuO[72] 69.1569.15[72] $52$ $66.9$ $7/2$ $3.3$
EuS[73] 16.5716.57[73] $13$ $16.7$ $7/2$ $0.8$
Fe2O3[74] 960960[74] $730$ $1022$ $5/2$ $6.4$
FeCl2[75] 23.5523.55[75] $11.5$ $23$ 11 $2.3$
FeF2[76] 78.478.4[76] $55$ $82.5$ 22 $5.2$
FeO[77] 198198[77] $148$ $222$ 22 $12.1$
FePS3[78] 120120[78] $91$ $136.5$ 22 $13.8$
FePSe3[79] 110110[79] $75$ $112.5$ 22 $2.3$
FePt3[80] 170170[80] $104$ $167.0$ $1.65$ $1.8$
HoFeO3[81] 647647[81] $456$ $638.4$ $5/2$ $1.3$
KCoF3[82] 114114[83] $33$ $99$ $1/2$ $13.2$
KCuF3[84] 3939[84] $15$ $45$ $1/2$ $15.4$
La2CoO4[85] 275275[85] $161$ $268.3$ $3/2$ $2.4$
La2NiO4[86] 328328[86] $220$ $440$ 11 $34.1$
LaFeO3[87] 738738[88] $504$ $705.6$ $5/2$ $4.4$
$T_{\mathrm{Exp.}}$(K)
$T_{\mathrm{MC}}$(K)
$T_{\mathrm{MC}}^{*}$(K)
S
Error(%)
LaMnO3[89] 139.5139.5[89] $101$ $151.5$ 22 $8.6$
LaVO3[87] 150150[90] $70$ $140$ 11 $6.7$
LiFePO4[91] 5050[91] $41$ $61.5$ 22 $23$
LuMnO3[92] 87.587.5[92] $71$ $106.5$ 22 $21.7$
MnBi[93] 630630[93] $394$ $591$ 22 $6.2$
MnBi2Te4[39] 2424[39] $16$ $22.4$ $5/2$ $6.7$
MnF2[20] 6767[20] $48$ $67.2$ $5/2$ $0.3$
MnO[94] 117117[94] $79$ $110.6$ $5/2$ $5.5$
MnPS3[95] 7878[95] $61$ $85.4$ $5/2$ $9.5$
MnPSe3[96] 7474[96] $52$ $72.8$ $5/2$ $1.6$
MnS[97] 152152[97] $105$ $147$ $5/2$ $3.3$
MnS2[98] 4848[98] $25$ $35$ $5/2$ $27.1$
MnTe[99] 310310[99] $251$ $351.4$ $5/2$ $13.4$
MnTiO3[100] 6565[100] $45$ $63$ $5/2$ $3.1$
MnWO4[101] 13.513.5[101] 99 $12.6$ $5/2$ $6.7$
NdFeO3[102] 693693[102] $463$ $648.2$ $5/2$ $6.5$
NiBr2[103] 5252[103] $28$ $56$ 11 $7.7$
NiCl2[104] 52.352.3[104] $36$ $72$ 11 $37.7$
NiF2[105] 73.273.2[105] $42$ $84$ 11 $14.8$
NiO[38] 523523[38] $310$ $620$ 11 $18.5$
PrFeO3[106] 707707[106] $471$ $659.4$ $5/2$ $6.7$
Rb2MnF4[107] 38.438.4[107] $33$ $46.2$ $5/2$ $20.3$
RbMnCl3[108] 9494[108] $69$ $96.6$ $5/2$ $2.8$
RbMnF3[109] 82.682.6[109] $54$ $75.6$ $5/2$ $8.5$
RbNiF3[110] 133133[110] $79$ $158$ 11 $18.8$
Sr2NiWO6[111] 5454[111] $26$ $52$ 11 $3.7$
SrMnBi2[112] 290290[112] $219$ $306.6$ $5/2$ $5.7$
SrMnSb2[113] 297297[114] $189$ $283.5$ 22 $4.5$
TbFeO3[115] 681681[115] $456$ $638.4$ $5/2$ $6.2$
TmFeO3[116] 635635[117] $442$ $618.8$ $5/2$ $2.6$
YbFeO3[118] 627627[119] $440$ $616$ $5/2$ $1.8$
YbMnBi2[17] 290290[17] $203$ $284.2$ $5/2$ $2.0$
YbMnSb2[16] 345345[16] $266$ $372.4$ $5/2$ $7.9$
YFeO3[120] 644.5644.5[121] $445$ $623$ $5/2$ $3.3$
YMnO3[122] 7070[122] $40$ $60$ 22 $14.3$
YVO3[123] 118118[123] $64$ $128$ 11 $8.5$